OpenStudy (anonymous):

what value of x will make the equation 5(6x + 8-2x) =-10

OpenStudy (texaschic101):

5(6x + 8 - 2x) = -10 5(4x + 8) = -10 4x + 8 = -10/5 4x + 8 = -2 4x = -2 - 8 4x = -10 x = -10/4 x = - 5/2 check.. 5(6x + 8 - 2x) = -10 5(4x + 8) = -10 5(4(-5/2) + 8) = -10 5(-10 + 8) = -10 5(-2) = -10 -10 = -10 (correct) if you replace x with -5/2, then the equation will be true

OpenStudy (texaschic101):

do you have any questions at all ? If so, please ask and I will explain the best I can